OpenStudy (dayton):

Help I need to solve for A 0=a+b+4 1= a((1+sqrt5)/2) + b((1-sqrt5)/2) + 8

OpenStudy (anonymous):

convert the first equation to be in terms of a or b

OpenStudy (dayton):

oh okay thank you

OpenStudy (anonymous):

come back and post your answer so we can see if its right

OpenStudy (dayton):

i end up with 1=a((1+sqrt5)/2)+(-a-4)((1-sqrt5)/2)+8 a comes out to be -5-2sqrt5 so b must be 9+2sqrt5

OpenStudy (anonymous):

i have it diferently i get \[\frac{ -10+3\sqrt{5} }{ 5 }\]

OpenStudy (anonymous):

as a and i get \[\frac{ -30+3\sqrt{5} }{ 5 }\] as b

OpenStudy (anonymous):

wait n sec let me start again i did that horribly wrong

OpenStudy (anonymous):

\[ 1=a(1+\frac{\sqrt5}{ 2 })+(-a-4)(1-\frac{\sqrt5}{ 2 })+8\] to \[1=a+\frac{ \sqrt{5} }{ 2 }a-a+\frac{ \sqrt{5} }{ 2 }a-4+2\sqrt{5}+8\]

OpenStudy (anonymous):

is that right?

OpenStudy (dayton):

so the answer should be -2-sqrt5 i think i just simplified wrong the first time

OpenStudy (anonymous):

\[1=\sqrt{5} a+4+2\sqrt{5}\]

OpenStudy (anonymous):

\[1-(4+2\sqrt{5})=\sqrt{5}a\]

OpenStudy (anonymous):

\[-3-2\sqrt{5}=\sqrt{5}a\]

OpenStudy (anonymous):

\[\frac{ -3-2\sqrt{5}}{ \sqrt{5} }=a\]

OpenStudy (anonymous):

we both seem to be wrong on this one

OpenStudy (anonymous):

what if if we do it as simultanoeus equations.....AO= a+b+41....(1) and AO=a(1+sqrt5)/2) +b(1-sqrt5)/2) +8....(2)